# Beyond the Black Box: Theory and Mechanism of Large Language Models

### v1

@@ 4.1 Fundamental Problems
These two questions, which concern the creation of foundational knowledge via pre-training and the adaptation of that knowledge via fine-tuning, form the theoretical bedrock of the Training Stage. A detailed illustration of the corresponding topics is shown in fig. 4. In what follows, we review the core theories and methods the community has developed to address these profound challenges.
@@ 4.2.1 Analysis on Pre-Training
The pre-training phase is where the model’s foundational capabilities are forged. This massive-scale, self-supervised process imbues the model with vast linguistic knowledge, factual information, and nascent reasoning abilities. The theoretical inquiry in this area focuses on some primary axes: understanding why and how the knowledge learned during pre-training is beneficial for downstream tasks, and formalizing the relationship between scale (data, parameters, compute) and capability, commonly known as Scaling Laws [kaplan2020scaling]. Building upon these, the community further seek for the explanation for the intelligence of LLM.
@@
The Benefits of Pre-Training. A significant body of theoretical work seeks to explain why self-supervised pre-training is so effective for transfer learning. Initial research provided a direct mathematical link, proving that a language model achieving $\epsilon$-optimal cross-entropy loss during pre-training can, in turn, enable a simple linear classifier to achieve an error rate of $\mathcal{O}(\sqrt{\epsilon})$ on downstream natural classification tasks [saunshi2020mathematical]. Subsequent work posits that pre-training enables the model to capture the underlying latent variable information within the text data [wei2021pretrained]. This analysis also helps explain the efficacy of different adaptation methods, noting that prompt-tuning requires weaker non-degeneration conditions than head-tuning, providing a theoretical basis for its strong performance in few-shot settings. The quality of the pre-training task itself is also a critical factor. zhao2023blessing provide a statistical theory demonstrating that high class diversity in the pre-training objective is key to improving the sample efficiency of downstream tasks.
@@
More general frameworks have been proposed to formalize the entire transfer process. deng2024generalization derives a generalization bound for the fine-tuned model, revealing that its performance depends on four key factors: Representation Transferrability, Representation-induced Rademacher Complexity, Domain Heterogeneity, and the generalization ability of the pre-training task itself. A unified perspective is further offered to theorize that pre-training learns the “contexture”—the top-$d$ singular functions of the association between inputs and their contexts [pmlr-v267-zhai25c]. A representation that successfully learns this contexture is proven to be optimal for downstream tasks that are compatible with that context. To avoid the costly process of fine-tuning to find the best model, zhang2025assessing introduce the DISCO framework. It uses Singular Value Decomposition (SVD) to analyze a model’s features, operating on the insight that different spectral components of the features have different degrees of transferability.
@@
Scaling Laws. Scaling laws are a set of empirical and theoretical principles that describe the predictable, power-law relationship between a model’s performance and increases in scale. A foundational work [hoffmann2022training] in this area establish that for compute-optimal training, model size ($N$) and the amount of training data ($D$) should be scaled proportionally with the compute budget ($C$), specifically $N_{opt}\propto C^{0.5}$ and $D_{opt}\propto C^{0.5}$. This finding reveals that many previous large-scale models were significantly undertrained. However, these laws can be disrupted. dohmatob2024tale provide a theoretical framework explaining that as synthetic, AI-generated data enters the training corpus, it can alter or break traditional scaling laws, leading to performance degradation and model collapse.
@@
Given the high cost of training, new methods for studying these laws have emerged. ruan2024observational propose a method that analyzes public models to bypass costly retraining, finding it can accurately predict complex performance changes, including phenomena previously considered “emergent”. The nature of emergence itself is also being explained by scaling. wu2025ushaped suggest emergence is not a mysterious qualitative shift but the result of two competing scaling patterns: difficult problems exhibit U-shaped scaling (getting worse before getting better), while simple problems show inverted-U scaling, with the “emergent” threshold appearing where these two trends interact.
@@
Deeper theoretical work seeks to explain why these power laws exist. bahri2024explaining identify four distinct scaling mechanisms: variance-limited and resolution-limited, each for both data and parameters. This work posits that the non-universal scaling exponents are linked to the intrinsic dimension of the data manifold. Further, havrilla2024understanding explicitly derive scaling exponents based on this manifold hypothesis.
@@
As the community looks to a data-constrained future, new scaling strategies are being explored. kim2025pre investigate the “data-constrained, compute-rich” regime, proposing a joint scaling recipe where both the number of ensemble members ($K$) and the parameters per member ($N$) are scaled to infinity. More recently, held2025relative move beyond absolute performance to study “relative” performance, showing that scaling is not a “universal equalizer”. The performance gaps between different data distributions evolve in different ways: some gaps converge (e.g., knowledge domains), while others diverge (e.g., certain AI risk behaviors).
@@
The Origin of Intelligence. Understanding the origins of intelligence in artificial neural networks remains a critical problem in AI research. Recently, compression has emerged as a popular perspective for understanding the success of Transformer models [Ilyatalk2023, LMisCompression, CompressionIntelligenceLinearly, pan2025understanding]. A prevailing view is that effective compression can give rise to intelligence [hutterprize, Ilyatalk2023, pan2025understanding]. Data compression focuses on removing redundant information, and from this perspective, Transformers can efficiently compress large-scale data while modeling the underlying target distribution using a limited number of parameters $\bm{\theta}$}. LMisCompression formalize the connection between the maximum likelihood training objective of LLMs and arithmetic coding, proposing that LLMs act as powerful lossless compressors. They also empirically demonstrate that foundation models can serve as general-purpose compressors. During training, a Transformer learns a parameterized distribution ${\bm{p}}_{{\bm{\theta}}}$ to maximize the log-likelihood, which is equivalent to minimizing the expected code length when the model is used for compression. According to Shannon’s source coding theorem [shannon], the minimum expected number of bits required to encode the data is precisely the entropy, which represents the theoretical limit of the model’s compression performance. ren2025revisiting further study Transformers in a controlled setting with a predefined target distribution, revealing an inherent bias toward learning distributions with lower entropy than the true target. This bias is primarily driven by the feed-forward (FFN) modules, highlighting a structural source of the model’s inductive preference.
@@
Beyond compression as a formal objective, it is believed to capture aspects of intelligence [hutter2005universal, Ilyatalk2023, pan2025understanding] to some extent. To quantify this, CompressionIntelligenceLinearly consider knowledge, commonsense, coding, and mathematical reasoning as proxies for intelligence and observe a strong linear relationship between compression efficiency and downstream task performance. pan2025understanding use the Kolmogorov Structure Function to show how models learn syntactic patterns first and factual knowledge according to frequency, connecting model capacity and data size to scaling laws. By linking compression efficiency to learning dynamics, these works provide a theoretical lens for understanding when LLMs generalize effectively versus when hallucinations occur, shedding light on the mechanisms underlying emergent intelligence.
@@ 4.2.2 Guarantees of Model Tuning
While pre-training forges the model’s foundational knowledge, fine-tuning is the critical process of adapting this general-purpose artifact to specific tasks or human intent [ouyang2022training, ren2025learning]. The advent of instruction tuning highlighted the necessity of this stage for achieving user alignment. This topic has catalyzed two major lines of theoretical inquiry. The first is the development and analysis of Parameter-Efficient Fine-Tuning (PEFT) methods, which seek to achieve adaptation by optimizing only a small subset of parameters, thereby dramatically reducing the computational burden [he2021towards, raffel2020exploring]. The second line of research delves into a more fundamental, mechanistic understanding of the adaptation process itself: how does fine-tuning alter the model’s internal computations? We will review these topics in the following parts.
@@
Parameter-Efficient Fine-Tuning. PEFT paradigms introduce a small number of new, learnable parameters while keeping the original model weights frozen. A key theoretical investigation [petrov2024when] provides insight into their expressive power. This work proves that while these methods are effective, they are less expressive than full fine-tuning. The role of attention in this process is further explored [oymak2023role], while subsequent improvements aim to enhance the mapping of input embeddings for better adaptation [tang2025adept].
@@
Nowadays, Low-Rank Adaptation (LoRA) [hu2022lora] has become a dominant PEFT strategy [mao2025survey]. The theoretical understanding of LoRA has advanced significantly. malladi2023kernel demonstrate that in the lazy regime, LoRA fine-tuning is nearly equivalent to full fine-tuning. As a theoretical foundation, zeng2024the analyze the expressive power of LoRA. From an optimization perspective, they show that LoRA can adapt any model $f$ to accurately represent any smaller target model $\tilde{f}$ if $\text{LoRA-rank}\geq(\text{width of }f)\times\frac{\text{depth of }\tilde{f}}{\text{depth of }f}$. jang2024lora have further proven that LoRA can eliminate spurious local minima, allowing gradient descent to find a high-performing low-rank solution. This is supported by another landscape analysis [pmlr-v267-kim25n], which shows that while other solutions exist, the standard zero-initialization and weight-decay mechanisms implicitly guide the optimization toward the desired low-rank global minimum.
@@
Deeper theoretical work has analyzed the individual components of LoRA. zhu2024asymmetry find an asymmetry in the learned matrices. This has led to an intense study of initialization strategies. While hayou2024impact suggests initializing $A$ randomly and $B$ with zeros allows for larger, more stable learning rates, pmlr-v267-li25bm challenge this, showing non-zero initialization can improve robustness to learning rate selection. A recent, theoretically-driven approach [pmlr-v267-zhang25ax] proves that LoRA adapters align with the singular subspace of the one-step full fine-tuning gradient. This insight leads to an initialization strategy based on this gradient, which is proven to converge linearly. The theoretical understanding has inspired new LoRA variants. hayou2024lora+ propose setting proportional learning rates for the $A$ and $B$ matrices. liu2024dora improve training stability by decomposing the pre-trained weights into magnitude and direction components, applying LoRA only to the direction component.
@@
An alternative line of research explores adaptation in other subspaces. bini2024ether use orthogonal transformations for fine-tuning. The theoretical connection between these methods and LoRA was then established [yuan2024bridging]. This concept of subspace training also underpins new optimizer-based PEFT methods. zhao2024galore propose a memory-efficient training strategy that performs gradient updates within a projected low-rank subspace. A follow-up work [pmlr-v267-he25i] analyzes the convergence properties and aim to guarantees convergence in typical stochastic settings. Other strategies include tuning only specific components, such as normalization layers, which has been shown to be surprisingly expressive [giannou2023expressive].
@@
Understanding Tuning Process. While PEFT methods offer practical recipes for adaptation, a fundamental theoretical question remains: how does the adaptation process actually alter the model’s internal computations and optimization landscape? A primary line of inquiry focuses on the optimization behavior of low-rank adapters. Unlike full-rank training, the introduction of low-rank constraints alters the loss landscape. liu2025on reveal that specific subspace optimization methods may possess superior optimization properties compared to standard LoRA. Furthermore, the training dynamics of LoRA itself exhibit distinct phases. Through a gradient flow perspective, pmlr-v258-xu25h identify that initialization scale is a critical factor. They theoretically prove that smaller initializations promote better alignment, thereby reducing the final error. On a broader scale, ijcai2025p760 attempts to provide a unified framework for selecting appropriate weight types and learning rates, offering theoretical guidance for the general fine-tuning of attention-based models.
@@
For methods that rely on modifying inputs or attention, theoretical analysis has focused on their expressive power and limitations. meyer2025memory formally prove a capacity bottleneck as the amount of information a Transformer can “memorize” via prompt tuning is linearly bounded by the prompt length. Furthermore, they demonstrate that for a single-layer Transformer, prompt tuning is restricted to generating outputs that lie within a specific hyperplane, highlighting significant expressive limitations compared to weight tuning. However, within these constraints, the attention mechanism plays a pivotal role. oymak2023role investigate the dynamics of soft prompts in a single-layer attention setting. They theoretically establish that softmax prompt attention is more expressive than self-attention or linear prompt attention in the context of mixture models. The study further characterizes how gradient descent naturally guides prompts to focus on sparse, task-relevant tokens. Additionally, pmlr-v267-diep25a establish a theoretical link between the “zero-initialized attention” mechanism and Mixture-of-Experts (MoE). They prove that this initialization strategy significantly improves sample efficiency compared to random initialization, with non-linear prompts theoretically outperforming linear ones.
@@
Finally, researchers are examining where the adaptation occurs and what structures are learned. Challenging the conventional wisdom that knowledge resides primarily in MLPs, he2025smt provide empirical and theoretical evidence that fine-tuning attention layers is more critical for downstream tasks than tuning MLP layers. This insight leads to the development of Sparse Matrix Tuning, which targets these high-impact parameters. Regarding the nature of the learned features, li2023transformers explore how fine-tuning affects semantic organization. They validate that this structural learning is a robust phenomenon that persists even when training is restricted to specific components. To further refine which components are adapted, jiang2025diffora introduce a differentiable adaptation matrix (DAM) to dynamically select modules for LoRA adaptation, theoretically proving that this selective approach can enhance convergence speed and generalization.

### v2

@@ 4.1 Fundamental Problems
These two questions, which concern the creation of foundational knowledge via pre-training and the adaptation of that knowledge via fine-tuning, form the theoretical bedrock of the Training Stage. A detailed illustration of the corresponding topics is shown in fig.˜4. In what follows, we review the core theories and methods the community has developed to address these profound challenges.
@@ 4.2.1 Analysis on Pre-Training
The pre-training phase is where the model’s foundational capabilities are forged. This massive-scale, self-supervised process imbues the model with vast linguistic knowledge, factual information, and nascent reasoning abilities. The theoretical inquiry in this area focuses on some primary axes: understanding why and how the knowledge learned during pre-training is beneficial for downstream tasks, and formalizing the relationship between scale (data, parameters, compute) and capability, commonly known as Scaling Laws (Kaplan et al., 2020a). Building upon these, the community further seek for the explanation for the intelligence of LLM.
@@
The Benefits of Pre-Training. A significant body of theoretical work seeks to explain why self-supervised pre-training is so effective for transfer learning. Initial research provided a direct mathematical link, proving that a language model achieving $\epsilon$-optimal cross-entropy loss during pre-training can, in turn, enable a simple linear classifier to achieve an error rate of $\mathcal{O}(\sqrt{\epsilon})$ on downstream natural classification tasks (Saunshi et al., 2020). Subsequent work posits that pre-training enables the model to capture the underlying latent variable information within the text data (Wei et al., 2021). This analysis also helps explain the efficacy of different adaptation methods, noting that prompt-tuning requires weaker non-degeneration conditions than head-tuning, providing a theoretical basis for its strong performance in few-shot settings. The quality of the pre-training task itself is also a critical factor. Zhao et al. (2023) provide a statistical theory demonstrating that high class diversity in the pre-training objective is key to improving the sample efficiency of downstream tasks.
@@
More general frameworks have been proposed to formalize the entire transfer process. Deng et al. (2024c) derive a generalization bound for the fine-tuned model, revealing that its performance depends on four key factors: Representation Transferrability, Representation-induced Rademacher Complexity, Domain Heterogeneity, and the generalization ability of the pre-training task itself. A unified perspective is further offered to theorize that pre-training learns the “contexture”—the top-$d$ singular functions of the association between inputs and their contexts (Zhai et al., 2025). A representation that successfully learns this contexture is proven to be optimal for downstream tasks that are compatible with that context. To avoid the costly process of fine-tuning to find the best model, Zhang et al. (2025c) introduce the DISCO framework. It uses Singular Value Decomposition (SVD) to analyze a model’s features, operating on the insight that different spectral components of the features have different degrees of transferability.
@@
Scaling Laws. Scaling laws are a set of empirical and theoretical principles that describe the predictable, power-law relationship between a model’s performance and increases in scale. A foundational work (Hoffmann et al., 2022b) in this area establish that for compute-optimal training, model size ($N$) and the amount of training data ($D$) should be scaled proportionally with the compute budget ($C$), specifically $N_{opt}\propto C^{0.5}$ and $D_{opt}\propto C^{0.5}$. This finding reveals that many previous large-scale models were significantly undertrained. However, these laws can be disrupted. Dohmatob et al. (2024) provide a theoretical framework explaining that as synthetic, AI-generated data enters the training corpus, it can alter or break traditional scaling laws, leading to performance degradation and model collapse.
@@
Given the high cost of training, new methods for studying these laws have emerged. Ruan et al. (2024) propose a method that analyzes public models to bypass costly retraining, finding it can accurately predict complex performance changes, including phenomena previously considered “emergent”. The nature of emergence itself is also being explained by scaling. Wu and Lo (2025) suggest emergence is not a mysterious qualitative shift but the result of two competing scaling patterns: difficult problems exhibit U-shaped scaling (getting worse before getting better), while simple problems show inverted-U scaling, with the “emergent” threshold appearing where these two trends interact.
@@
Deeper theoretical work seeks to explain why these power laws exist. Bahri et al. (2024) identify four distinct scaling mechanisms: variance-limited and resolution-limited, each for both data and parameters. This work posits that the non-universal scaling exponents are linked to the intrinsic dimension of the data manifold. Further, Havrilla and Liao (2024) explicitly derive scaling exponents based on this manifold hypothesis.
@@
As the community looks to a data-constrained future, new scaling strategies are being explored. Kim et al. (2025d) investigate the “data-constrained, compute-rich” regime, proposing a joint scaling recipe where both the number of ensemble members ($K$) and the parameters per member ($N$) are scaled to infinity. Held et al. (2025a) move beyond absolute performance to study “relative” performance, showing that scaling is not a “universal equalizer”. The performance gaps between different data distributions evolve in different ways: some gaps converge (e.g., knowledge domains), while others diverge (e.g., certain AI risk behaviors). More recently, Liu et al. (2025h) introduce the concept of superposition, where models represent more features than the dimensions they have, and thus propose a power-law loss with model size without assuming power laws in data. Zou et al. (2026a) further introduce the concept of effective frontiers and unify existing scaling laws.
@@
The Origin of Intelligence. Understanding the origins of intelligence in artificial neural networks remains a critical problem in AI research. Recently, compression has emerged as a popular perspective for understanding the success of Transformer models (Sutskever, 2023; Delétang et al., 2023; Huang et al., 2024e; Pan et al., 2025c). A prevailing view is that effective compression can give rise to intelligence (Hutter, 2006; Sutskever, 2023; Pan et al., 2025c). Data compression focuses on removing redundant information, and from this perspective, Transformers can efficiently compress large-scale data while modeling the underlying target distribution using a limited number of parameters $\bm{\theta}$}. Delétang et al. (2023) formalize the connection between the maximum likelihood training objective of LLMs and arithmetic coding, proposing that LLMs act as powerful lossless compressors. They also empirically demonstrate that foundation models can serve as general-purpose compressors. During training, a Transformer learns a parameterized distribution ${\bm{p}}_{{\bm{\theta}}}$ to maximize the log-likelihood, which is equivalent to minimizing the expected code length when the model is used for compression. According to Shannon’s source coding theorem (Shannon, 1948), the minimum expected number of bits required to encode the data is precisely the entropy, which represents the theoretical limit of the model’s compression performance. Ren and Liu (2025) further study Transformers in a controlled setting with a predefined target distribution, revealing an inherent bias toward learning distributions with lower entropy than the true target. This bias is primarily driven by the feed-forward (FFN) modules, highlighting a structural source of the model’s inductive preference.
@@
Beyond compression as a formal objective, it is believed to capture aspects of intelligence (Hutter, 2005; Sutskever, 2023; Pan et al., 2025c) to some extent. To quantify this, Huang et al. (2024e) consider knowledge, commonsense, coding, and mathematical reasoning as proxies for intelligence and observe a strong linear relationship between compression efficiency and downstream task performance. Pan et al. (2025c) use the Kolmogorov Structure Function to show how models learn syntactic patterns first and factual knowledge according to frequency, connecting model capacity and data size to scaling laws. By linking compression efficiency to learning dynamics, these works provide a theoretical lens for understanding when LLMs generalize effectively versus when hallucinations occur, shedding light on the mechanisms underlying emergent intelligence.
@@ 4.2.2 Guarantees of Model Tuning
While pre-training forges the model’s foundational knowledge, fine-tuning is the critical process of adapting this general-purpose artifact to specific tasks or human intent (Ouyang et al., 2022; Ren and Sutherland, 2025). The advent of instruction tuning highlighted the necessity of this stage for achieving user alignment. This topic has catalyzed two major lines of theoretical inquiry. The first is the development and analysis of Parameter-Efficient Fine-Tuning (PEFT) methods, which seek to achieve adaptation by optimizing only a small subset of parameters, thereby dramatically reducing the computational burden (He et al., 2021; Raffel et al., 2020b). The second line of research delves into a more fundamental, mechanistic understanding of the adaptation process itself: how does fine-tuning alter the model’s internal computations? We will review these topics in the following parts.
@@
Parameter-Efficient Fine-Tuning. PEFT paradigms introduce a small number of new, learnable parameters while keeping the original model weights frozen. A key theoretical investigation (Petrov et al., 2024) provides insight into their expressive power. This work proves that while these methods are effective, they are less expressive than full fine-tuning. The role of attention in this process is further explored (Oymak et al., 2023), while subsequent improvements aim to enhance the mapping of input embeddings for better adaptation (Tang et al., 2025a).
@@
Nowadays, Low-Rank Adaptation (LoRA) (Hu et al., 2022) has become a dominant PEFT strategy (Mao et al., 2025). The theoretical understanding of LoRA has advanced significantly. Malladi et al. (2023) demonstrate that in the lazy regime, LoRA fine-tuning is nearly equivalent to full fine-tuning. As a theoretical foundation, Zeng and Lee (2024) analyze the expressive power of LoRA. From an optimization perspective, they show that LoRA can adapt any model $f$ to accurately represent any smaller target model $\tilde{f}$ if $\text{LoRA-rank}\geq(\text{width of }f)\times\frac{\text{depth of }\tilde{f}}{\text{depth of }f}$. Jang et al. (2024) have further proven that LoRA can eliminate spurious local minima, allowing gradient descent to find a high-performing low-rank solution. This is supported by another landscape analysis (Kim et al., 2025c), which shows that while other solutions exist, the standard zero-initialization and weight-decay mechanisms implicitly guide the optimization toward the desired low-rank global minimum.
@@
Deeper theoretical work has analyzed the individual components of LoRA. Zhu et al. (2024c) find an asymmetry in the learned matrices. This has led to an intense study of initialization strategies. While Hayou et al. (2024b) suggest initializing $A$ randomly and $B$ with zeros allows for larger, more stable learning rates, Li et al. (2025d) challenge this, showing non-zero initialization can improve robustness to learning rate selection. A recent, theoretically-driven approach (Zhang et al., 2025d) proves that LoRA adapters align with the singular subspace of the one-step full fine-tuning gradient. This insight leads to an initialization strategy based on this gradient, which is proven to converge linearly. The theoretical understanding has inspired new LoRA variants. Hayou et al. (2024a) propose setting proportional learning rates for the $A$ and $B$ matrices. Liu et al. (2024c) improve training stability by decomposing the pre-trained weights into magnitude and direction components, applying LoRA only to the direction component.
@@
An alternative line of research explores adaptation in other subspaces. Bini et al. (2024) use orthogonal transformations for fine-tuning. The theoretical connection between these methods and LoRA was then established (Yuan et al., 2024). This concept of subspace training also underpins new optimizer-based PEFT methods. Zhao et al. (2024b) propose a memory-efficient training strategy that performs gradient updates within a projected low-rank subspace. A follow-up work (He et al., 2025b) analyzes the convergence properties and aim to guarantees convergence in typical stochastic settings. Other strategies include tuning only specific components, such as normalization layers, which has been shown to be surprisingly expressive (Giannou et al., 2023a).
@@
Understanding Tuning Process. While PEFT methods offer practical recipes for adaptation, a fundamental theoretical question remains: how does the adaptation process actually alter the model’s internal computations and optimization landscape? A primary line of inquiry focuses on the optimization behavior of low-rank adapters. Unlike full-rank training, the introduction of low-rank constraints alters the loss landscape. Liu et al. (2025f) reveal that specific subspace optimization methods may possess superior optimization properties compared to standard LoRA. Furthermore, the training dynamics of LoRA itself exhibit distinct phases. Through a gradient flow perspective, Xu et al. (2025c) identify that initialization scale is a critical factor. They theoretically prove that smaller initializations promote better alignment, thereby reducing the final error. On a broader scale, Yao et al. (2025c) attempts to provide a unified framework for selecting appropriate weight types and learning rates, offering theoretical guidance for the general fine-tuning of attention-based models.
@@
For methods that rely on modifying inputs or attention, theoretical analysis has focused on their expressive power and limitations. Meyer et al. (2025) formally prove a capacity bottleneck as the amount of information a Transformer can “memorize” via prompt tuning is linearly bounded by the prompt length. Furthermore, they demonstrate that for a single-layer Transformer, prompt tuning is restricted to generating outputs that lie within a specific hyperplane, highlighting significant expressive limitations compared to weight tuning. However, within these constraints, the attention mechanism plays a pivotal role. Oymak et al. (2023) investigate the dynamics of soft prompts in a single-layer attention setting. They theoretically establish that softmax prompt attention is more expressive than self-attention or linear prompt attention in the context of mixture models. The study further characterizes how gradient descent naturally guides prompts to focus on sparse, task-relevant tokens. Additionally, Diep et al. (2025) establish a theoretical link between the “zero-initialized attention” mechanism and Mixture-of-Experts (MoE). They prove that this initialization strategy significantly improves sample efficiency compared to random initialization, with non-linear prompts theoretically outperforming linear ones.
@@
Finally, researchers are examining where the adaptation occurs and what structures are learned. Challenging the conventional wisdom that knowledge resides primarily in MLPs, He et al. (2025a) provide empirical and theoretical evidence that fine-tuning attention layers is more critical for downstream tasks than tuning MLP layers. This insight leads to the development of Sparse Matrix Tuning, which targets these high-impact parameters. Regarding the nature of the learned features, Li et al. (2023c) explore how fine-tuning affects semantic organization. They validate that this structural learning is a robust phenomenon that persists even when training is restricted to specific components. To further refine which components are adapted, Jiang et al. (2025) introduce a differentiable adaptation matrix (DAM) to dynamically select modules for LoRA adaptation, theoretically proving that this selective approach can enhance convergence speed and generalization.